OpenStudy (anonymous):

how do you write y^2/5 as a radical

sam (.sam.):

\[\sqrt[5]{y ^{2}}\]

sam (.sam.):

\[\huge y^{\frac{a}{b}}=\sqrt[b]{y ^{a}}\]

OpenStudy (muhammad_nauman_umair):

\[\sqrt[5]{y^2}\]